$32,000 $24,500 $20,000 $15,313 During its first year of operations, Silverman Company paid $10,000 for direct...

$32,000

$24,500

$20,000

$15,313

During its first year of operations, Silverman Company paid $10,000 for direct materials and $11,500 for production workers' wages. Lease payments and utilities on the production facilities amounted to $10,500 while general, selling, and administrative expenses totaled $3,000. The company produced 8,000 units and sold 5,000 units at a price of $6.50 a unit.

What is Silverman's cost of goods sold for the year?
Answer #1

Cost of production = Direct materials + Direct labour + Overheads

= 10,000 + 11,500 + 10,500

= 32,000

Cost of production per unit = 32,000 / 8,000 = 4

Cost of goods sold = 5,000 units * 4 per uniit = 20,000